# NeuroLGP-SM: Scalable Surrogate-Assisted Neuroevolution for Deep Neural Networks

### v1

@@ I Introduction
Evolutionary Algorithms (EAs) [9] have proven to be effective in both the crafting of architectures and hyperparameter optimisation of Deep Neural Networks (DDNs) [25]. This application is commonly known as neuroevolution, a widely explored field highlighted by the abundance of scientific publications and impactful outcomes [11], and have been applied to numerous problem domains, such as autonomous vehicles [14], face recognition [8]. The pursuit of optimal DNN architectures has led to various methodologies, including EAs [11], reinforcement learning [44], etc. However, a significant challenge persists across these methods: the substantial computational resources required to identify high-performing networks.
The rise of GPU-accelerated hardware has helped alleviate some of this computational cost, however, a significant proportion of research in DNNs is based on incremental improvements on DNN algorithms for benchmark problems [33], where there is a significant correlation between network complexity for incremental gains in terms of additional performance. In fact, when looking at very large models of hundreds of billions of parameters, it can cost millions of dollars for a single iteration [26]. This energy consumption is further compounded when considering population-based neuroevolutionary techniques which require many networks to be trained and evaluated in order to find suitable architectures.
@@ II-A Neuroevolution using Surrogate Models
Freeze-Thaw Bayesian Optimisation (FBO), proposed by Swersky et al.  [43], uses Bayesian optimisation to determine whether a particular neural network that has been partially trained should be fully evaluated. This approach is novel in that the system stops training (or freezing) of less promising networks, instead spending valuable resources on the most promising networks. Of note, is the fact that the FBO approach relies on the phenotypic behaviour of DNNs. It is important to note the majority of works rely on genotypic information when building surrogate models for neuroevolution [10, 41, 17]. For reference, a major work in this regard is that of Sun et al. [41] referred to as the End-to-End Performance Predictor. This approach uses an offline surrogate model based on random forests to accelerate learning of CNN architectures. The CNN architecture is encoded such that it maps to a numerical decision variable, which is passed to the random forest-based surrogate-model. Not only does this approach not require large amounts of training time but, also alleviates the requirement of a smooth learning curve [43].
@@ III-A Surrogate assisted models: Kriging and Kriging Partial Least Squares
Kriging is an interpolation-based technique that assumes spatial correlation exists between known data points, based on the distance, and variation between these points. We aim to use observations $\{y(x_{1}),y(x_{2}),...y(x_{n})\}$ to help estimate an unknown function value $\hat{y}(x^{{}^{\prime}})$ for the unknown data point $x^{{}^{\prime}}$, where $n$ is the number of individuals in the training data. A kernel function $K(\cdot)$ is used to express the spatial correlation between two samples $x_{i}$ and $x^{{}^{\prime}}_{i}$ as shown in Eq. 1,

### v3

@@ I Introduction
Evolutionary Algorithms (EAs) [10] have proven to be effective in both the crafting of architectures and hyperparameter optimisation of Deep Neural Networks (DDNs) [25]. This application is commonly known as neuroevolution [11], and have been applied to numerous problem domains, such as autonomous vehicles [14, 39] and face recognition [9, 8]. The pursuit of optimal DNN architectures has led to various methodologies, including EAs [11], reinforcement learning [44], etc. However, a significant challenge persists across these methods: the substantial computational resources required to identify high-performing networks.
The rise of GPU-accelerated hardware has helped alleviate some of this computational cost, however, a significant proportion of research in DNNs is based on incremental improvements on DNN algorithms for benchmark problems, where there is a significant correlation between network complexity for incremental gains in terms of additional performance. In fact, when looking at very large models of hundreds of billions of parameters, it can cost millions of dollars for a single iteration [26]. This energy consumption is further compounded when considering population-based neuroevolutionary techniques which require many networks to be trained and evaluated in order to find suitable architectures.
@@ II-A Neuroevolution using Surrogate Models
Freeze-Thaw Bayesian Optimisation (FBO), proposed by Swersky et al. [43], uses Bayesian optimisation to determine whether a particular neural network that has been partially trained should be fully evaluated. This approach is novel in that the system stops training (or freezing) of less promising networks, instead spending valuable resources on the most promising networks. Of note, is the fact that the FBO approach relies on the phenotypic behaviour of DNNs. It is important to note the majority of works rely on genotypic information when building surrogate models for neuroevolution [41, 17]. For reference, a major work in this regard is that of Sun et al. [41] referred to as the End-to-End Performance Predictor. This approach uses an offline surrogate model based on random forests to accelerate learning of CNN architectures. The CNN architecture is encoded such that it maps to a numerical decision variable, which is passed to the random forest-based surrogate-model. Not only does this approach not require large amounts of training time but, also alleviates the requirement of a smooth learning curve [43].
@@ III-A Surrogate assisted models: Kriging and Kriging Partial Least Squares
Kriging is an interpolation-based technique that assumes spatial correlation exists between known data points, based on the distance, and variation between these points. We aim to use observations $\{y(x_{1}),y(x_{2}),...y(x_{n})\}$ to help estimate an unknown function value $\hat{y}(x^{*})$ for the unknown data point $x^{*}$, where $n$ is the number of individuals in the training data. A kernel function $K(\cdot)$ is used to express the spatial correlation between two samples $x_{i}$ and $x^{{}^{\prime}}_{i}$ as shown in Eq. 1,